The length of a rectangle is 2 more than three times the width. If the perimeter is 92 cm, find the dimensions of the rectangle.
Answer by anantha(86)

sol:
let the width be x
three times the fwidth=3x
length=2 more than 3 times the width
length=2+3x
perimeter of the rectangle=2(length+width)
=2(2+3x+x)
=2(2+4x)
=2*2+2*4x
=4+8x
according to the problem perimeter of the rectangle=92cm
4+8x=92
subtract 4 on both sides
4+8x-4=92-4
8x=88
dividing by 8 on both sides
8x/8=88/8
x=11cm
width=11cm
length=2+3x
=2+3*11
=2+33
=35
length=35cm